# Localized kernel method for separation of linear chirps

### v3

@@ 1 Introduction
The task of separating a superposition of signals into its individual components is a common challenge encountered in various signal processing applications, especially in domains such as audio [25, 28, 38] and radar [18, 11, 8, 3, 29, 23]. Due to the extensive set of applications and the frequent occurrence of signal source separation challenges, methodologies are often tailored to the specific field in question, as well as the level of prior knowledge of both the component signals and the mixing mechanism.
@@
The instantaneous frequency of the chirp $s_{n}$ in (1.2) at time $t$ is defined by $\omega+(2B/d)(t-n*\mbox{PRI})$. Our approach is to assume (following [5, 4]) that on a sufficiently small interval of the form $t-\Delta,t+\Delta]$, this frequency is approximated by a constant frequency at the center of the interval. We will then use specially constructed localized kernels to solve the problem for each of these intervals as described in [26, 4]. A more elaborate subdivision is necessary when the instantaneous frequencies of different pulses cross each other. We will demonstrate experimentally that our method is very stable, capable of working with a very small SNR if the sampling rate is sufficiently high. Thus, the only limitation is the hardware capability, not a theoretical limitation. We will analyze theoretically the connection between $\Delta$ and the various parameters of the pulses and the accuracy.
@@
We will describe some related work in Section 2. The mathematical foundation behind our work is explained in Section 3, including the intuition and definition of the Signal Separation Operator (SSO) introduced in [4]. The algorithms and methodology to use SSO for the chirp separation problem are described in Section 4 with step-by-step illustrations on one example. In particular, we demonstrate the connection between the sampling rate, noise level, minimal separation among the instantaneous frequencies. We have tested our algorithm on a database of 7 examples. The process to generate this data is described in Section 5. Section 6 describes some heuristics on the choice of the tunable parameters listed in Table 1. A full report on the entire dataset which we worked on is discussed in Section 7, and the results summarized in the appendix.
@@ 2 Related works
In the case when the frequencies are constant, the problem is known as parameter estimation in exponential sums. This problem apparently goes back to the 1795 paper by Prony [7]. Mathematically, the problem can be stated as follows. Let $K\geq 1$ be an integer, $\{\lambda_{j}\}_{k=1}^{K}\subset(-\pi,\pi]$, $\mu=\sum_{j=1}^{K}A_{j}\delta_{\lambda_{j}}$ for some complex numbers $\{A_{j}\}_{k=1}^{K}$, where $\delta_{x}$ denotes the Dirac delta supported at $x$. Given noisy variants of the Fourier coefficients
@@
for some integer $N\geq 1$ (where $\hat{\mathbf{i}}=\sqrt{-1}$), determine the locations $\lambda_{j}$ and the coefficients $A_{j}$. Of course, finding $\lambda_{j}$’s is a bigger challenge; once we obtain these, the $A_{j}$’s can be found using a least square fit for the data (2.1). We will not go into a survey of the very large amount of literature on this problem, except to point out that our current work is based on ideas developed in [26, 4, 27].
If the radar pulses have sufficiently high SNR, then they can be detected using an energy-based envelope detector operating on the complex valued time-series data, known as In-phase and Quadrature (IQ) data [20, 19, 12]. It is actually more common for modern solid state radars to transmit low probability of intercept (LPI) waveforms that utilize LFM pulses at low and negative SNR [37, 30]; in this case, the envelope detector will no longer be suitable. In the low power setting, there are two broad classes of methods, the first operates on the IQ data directly, while the second utilizes time-frequency transforms approaching detection and separation as a 2D signal processing problem. In the time domain, there are analytic methods that operate on the IQ data directly, such as variations of independent components algorithm (ICA) [18, 22, 16, 33], with extensions to nonlinear chirps [32, 39], and incorporation of simple neural networks [31]. A more common approach is to utilize time-frequency transforms and process the result as an image instead of operating on the IQ data directly [17, 35, 2, 14, 24, 34].
@@
The quantity $\phi_{j}^{\prime}(t)$ is known as the instantaneous frequency (IF) of the component $f_{j}$. The problem is to estimate the IFs and the IAs. This is also a very important and old problem, starting with a paper by a Nobel Laureate, Gabor [13]. One of the first approaches to solve the problem is the Hilbert-transform based empirical mode decomposition (EMD) by Huang [15]. However, there is no mathematical theory behind this approach. A mathematically rigorous approach was given by Daubeschies and her collaborators [6, 5]. These seminal papers have given rise to a lot of variations and improvements. A complete survey is both beyond the scope of this paper as well as not relevant here. We refer to [21] for more references on this topic. In our paper, we will use the algorithm given in [36] for comparing our method with SST.
In [4], we have applied the technique based on localized trigonometric kernels to define a signal separation operator (SSO) to solve this problem under conditions weaker than those in [5]. Our conditions do assume that all the components are present in the interval $[0,T]$; i.e., the number of signals $f_{j}$ is the same throughout the interval, and there is no discontinuity in any of these. Our method finds the IFs by looking at the peaks of a power spectrum rather than an elaborate reference curve mechanism of SST. Likewise, the IAs are estimated by a simple substitution compared to the limiting process required in SST. Finally, our method can be implemented very fast using FFT.
The paper [21] deals specifically with the separation of chirplets, based on a modification of a filter described in [4]. It is not clear when the various conditions in this paper are satisfied. Also, the method requires a three dimensional search. In this paper, we will demonstrate how the methods developed in [4] can be adapted for the chirp separation problem, including the case of crossover frequencies.
@@ 3 Mathematical foundations
In this section, we start in Section 3.1 by explaining the basic idea and theorem behind our methods in the case of signals with constant amplitudes and frequencies. In Section 3.2, we explain how the general problem of blind source signal separation can be reduced to this case at each point under certain assumptions as in [5, 4]. A theoretical analysis of the performance of SSO in the presence of noise is given in Section 3.3.
@@ 3.1 Constant amplitudes and frequencies
where $\epsilon_{\ell}$’s are independent realizations of a sub-Gaussian random variable. This discussion summarizes the corresponding results in [27]. Following [1, Section 2.3], we recall that a real-valued random variable $X$ is called sub-Gaussian with parameter $V$ ($X\in\mathcal{G}(V)$) if $\mathbb{E}(X)=0$ and $\log\mathbb{E}(\exp(tX))\leq tV^{2}/2$ for all $t>0$; e.g., if $X$ is a mean zero normal random variable with variance $V$, then $X\in\mathcal{G}(V)$. Likewise, any bounded random variable is sub-Gaussian. We define a complex valued random variable $X$ to be in $\mathcal{G}(V)$ if its real and imaginary parts are in $\mathcal{G}(V)$. It is not difficult to see that if $X_{1},\cdots,X_{n}$ are i.i.d., complex valued variables all in $\mathcal{G}(V)$, $\mathbf{a}=(a_{1},\cdots,a_{n})\in{\mathbb{C}}^{n}$, $|\mathbf{a}|_{n}^{2}=\sum_{\ell=1}^{n}|a_{\ell}|^{2}$, then $\sum_{\ell=1}^{n}a_{\ell}X_{\ell}\in\mathcal{G}(|\mathbf{a}|_{n}V)$, and hence, that
@@
Explicit expressions for $L$ in terms of $H$ and $S$ are given in [10]. The estimate (3.7) implies that $\hbar_{n}^{-1}\Phi_{n}(x)$ is an approximation to the Dirac delta $\delta_{0}$.
@@
Because of the localization of $\Phi_{n}$, one can deduce that $\sigma_{n}(\mu)(x)$ is small when $x$ is away from each of the $\lambda_{j}$’s, and when it is large then $x$ must be close to exactly one of the $\lambda_{j}$’s. Moreover, $E_{n}(x)$ being a weighted average of the noise realizations is small for large enough $n$. These observations are made precise in the following Theorem 3.1, proved in [27].
@@
Before stating the main theorem, we estimate the noise term $E_{n}(x)$ in the following lemma [27, Lemma 4.1].
@@ 3.2 Reduction to the constant parameter case
The following theorem is a reformulation of [4, Theorem 4.2]. We will reproduce the proof for the sake of completeness.
@@ 3.3 Signal separation operator
This leads to the following definition (cf. [4, Definition 2.3])
@@
where for each $t$, $\epsilon(t)$ is a realization of a complex sub-Gaussian random variable in $\mathcal{G}(V)$. Theorem 3.1 can be translated into Theorem 3.3 below, where we abuse the notation to make the comparison easier. We note that Theorem 3.3 itself is a restatement of [4, Theorems 2.4,4.1] in a somewhat modified form taking the noise into account. In order to state this theorem, we abuse the notation in Section 3.1 again, and write (in addition to (3.19))
@@ 4 Algorithms
The clusters defined in Theorem 3.3 are found, using a clustering algorithm. For example, in this paper we used DBSCAN [9]. The notation we use in our algorithm is
@@
Algorithm 1 is designed to find the IFs. A challenge here is to determine the correct threshold related to the unknown minimal amplitude $\mathfrak{m}$ in the presence of noise. A value which is too small would indicate superfluous frequencies, a value which is too large would miss some frequencies. This effect is discussed in our paper [27]. There are other details such as detecting the clusters and find the right peaks which are described below. Part 2 will estimate the parameters of the instantaneous frequencies obtained by part 1 using least squares. The challenge here is to figure out where the different signal components begin and end, and which part of the signal to use in order to set up the least square problem in order to obtain reliable solutions. SSO does not work directly when there are crossover frequencies. In fact, the set up itself does not make sense in general in such cases; e.g., whether the crossover of the form $X$ represents two linear frequencies crossing or two $V$-shaped frequencies touching each other. Part 3 is the refinement process where we use the facts that SSO is inherently unable to resolve crossovers and that we are dealing with linear chirps to detect the presence of such frequencies. If sufficiently long linear chirps are present on both sides of the crossover, then we can detect those using the previous two algorithms with a smaller value of $\Delta$ and interpolate to determine where the crossover happens and which chirps cross each other.
@@ 7.3 Comparison with SST
We recall that the most often used method for finding IF’s is EMD, which is purely heuristic. The method known as the synchrosqueezing transform method (SST) is based on a solid mathematical foundation and works better than EMD. In this section, we compare our results with those obtained by using SST. As mentioned earlier, there are many versions of SST. We use the implementation of the SST algorithm as described in Section III of [36].
In the experiment, we use the code from GitHub repository (https://github.com/ebrevdo/synchrosqueezing) from [36]. We use the same parameters used by the authors in their examples, provided in the code, which use the shifted Gaussian (or Morlet) wavelet with zero mean. We refer the reader to [36] for a more detailed discussion of the approach and parameters.
@@ 8 Conclusions
Separation of a superposition of blind source signals is an important problem in audio and radar signal processing, with many applications, including electronic warfare and electronic intelligence. In this paper, we have modified a method proposed in [4] for solving this problem in the case when the instantaneous frequencies of the components are linear. The method is based on a filtered FFT and hence is very fast. While the paper [4] assumes the frequencies to be slowly and continuously varying, our modification includes the case when the signals are discontinuous and when there are crossover frequencies present. We analyze theoretically the behavior of SSO in the presence of noise, giving insight into the relationship between the sampling frequency, desired accuracy, and noise. We have illustrated the effectiveness of our method with a few well chosen examples and tested it on a small database of 7 signals, demonstrating consistent behavior. Our method outperforms the SST algorithm given in [36].

### v4

@@ 1 Introduction
The task of separating a superposition of signals into its individual components is a common challenge encountered in various signal processing applications, especially in domains such as audio [makino2018, Nugraha2016, yuan2025] and radar [Joshil2020, Filice2013, Dutt2019, Wang2022, Nuhoglu2023, Yanchao2018]. Due to the extensive set of applications and the frequent occurrence of signal source separation challenges, methodologies are often tailored to the specific field in question, as well as the level of prior knowledge of both the component signals and the mixing mechanism.
@@
The instantaneous frequency of the chirp $s_{n}$ in (1.2) at time $t$ is defined by $\omega+(2B/d)(t-n*\mbox{PRI})$. Our approach is to assume (following [ingrid2011, bspaper]) that on a sufficiently small interval of the form $t-\Delta,t+\Delta]$, this frequency is approximated by a constant frequency at the center of the interval. We will then use specially constructed localized kernels to solve the problem for each of these intervals as described in [loctrigwave, bspaper]. A more elaborate subdivision is necessary when the instantaneous frequencies of different pulses cross each other. We will demonstrate experimentally that our method is very stable, capable of working with a very small SNR if the sampling rate is sufficiently high. Thus, the only limitation is the hardware capability, not a theoretical limitation. We will analyze theoretically the connection between $\Delta$ and the various parameters of the pulses and the accuracy.
Even though the problem is formulated in terms of a continuous variable, we only have discrete samples of the form $F(\ell\delta)$, $|\ell|<N$. Therefore, we cannot use techniques based on the Fourier transform, including short term Fourier transform. This would require a discretization, resulting in many errors: discretization error, aliasing, etc. One could think of extending $F$ as a $2N\delta$ periodic function, and using harmonic analysis on the group ${\mathbb{Z}}_{N}$. This would restrict the instantaneous frequencies to be lattice points.
Our approach is simpler. First, we will show in Theorem 3.2 that for any sufficiently small interval of the form $t^{*}-\Delta,t^{*}+\Delta]$, $F(\ell\delta)$ can be approximated by a quantity of the form $\tilde{\mu}(\ell)=\sum_{k=1}^{K(t^{*})}A_{k}(t^{*})\exp(-\hat{\mathbf{i}}\lambda_{k}(t^{*})\ell)+\mbox{noise}$, where $\lambda_{k}(t^{*})$’s are (unitless) instantaneous frequencies at $t^{*}$, scaled by the sampling parameter $\delta$. For a sufficiently large sampling rate (corresponding to a small $\delta$), we can ensure that each $\lambda_{k}(t^{*})\in[0,\pi]$.
We then revert our thinking about the physical notions of time and frequency. The points $t^{*}$ and the end points $t^{*}\pm\Delta$ need to be chosen so as to be among the grid points $\ell\delta$ at which the signal is sampled. In our reverse thinking, we now identify the sampling point $\ell\delta$ in the interval $t^{*}-\Delta,t^{*}+\Delta]$ with the “frequency” $\ell$ in the sense of trigonometric Fourier series. This has the advantage that we can use the samples directly without having to go through any transformations to convert them into the “Fourier domain”. In particular, this bypasses problems like discretization, aliasing, circularity, etc. which would arise if we thought in terms of Fourier transform.
The instantaneous frequencies $\lambda_{k}(t^{*})$ are now viewed as unknown points in the “time domain” ${\mathbb{R}}/(2\pi{\mathbb{Z}})$ to be determined. In particular, we no longer need to require $\lambda_{k}(t^{*})$’s to be grid points.
[…]
$\mu(x)=\sum_{\ell\in{\mathbb{Z}}}\tilde{\mu}(\ell)\exp(i\ell x)=\sum_{\ell\in{\mathbb{Z}}}F(\ell\delta)\exp(i\ell x),\qquad x\in{\mathbb{R}}/(2\pi{\mathbb{Z}}).$
Even then, this Fourier series will not converge, and in any case, we have the values of $F(\ell\delta)$ only for $|\ell|<N$ for some $N$. In the classical theory of trigonometric approximation, certain summability factors are introduced to obtain an approximation of a function given finitely many of its Fourier coefficients. We have constructed a highly localized summability kernel for this purpose. This enables us both to give quantitative estimates on the error in the estimation of the instantaneous frequencies, as well achieve a substantial noise reduction.
@@
We will describe some related work in Section 2. The mathematical foundation behind our work is explained in Section 3, including the intuition and definition of the Signal Separation Operator (SSO) introduced in [bspaper]. The algorithms and methodology to use SSO for the chirp separation problem are described in Section 4 with step-by-step illustrations on one example. In particular, we demonstrate the connection between the sampling rate, noise level, minimal separation among the instantaneous frequencies. We have tested our algorithm on a database of 7 examples. The process to generate this data is described in Section 5. Section 6 describes some heuristics on the choice of the tunable parameters listed in Table 1. A full report on the entire dataset which we worked on is discussed in Section 7, and the results summarized in the appendix.
@@ 2 Related works
In the case when the frequencies are constant, the problem is known as parameter estimation in exponential sums. This problem apparently goes back to the 1795 paper by Prony [prony_original]. Mathematically, the problem can be stated as follows. Let $K\geq 1$ be an integer, $\{\lambda_{j}\}_{k=1}^{K}\subset(-\pi,\pi]$, $\mu=\sum_{j=1}^{K}A_{j}\delta_{\lambda_{j}}$ for some complex numbers $\{A_{j}\}_{k=1}^{K}$, where $\delta_{x}$ denotes the Dirac delta supported at $x$. Given noisy variants of the Fourier coefficients
@@
for some integer $N\geq 1$ (where $\hat{\mathbf{i}}=\sqrt{-1}$), determine the locations $\lambda_{j}$ and the coefficients $A_{j}$. Of course, finding $\lambda_{j}$’s is a bigger challenge; once we obtain these, the $A_{j}$’s can be found using a least square fit for the data (2.1). We will not go into a survey of the very large amount of literature on this problem, except to point out that our current work is based on ideas developed in [loctrigwave, bspaper, mhaskar2024robust].
If the radar pulses have sufficiently high SNR, then they can be detected using an energy-based envelope detector operating on the complex valued time-series data, known as In-phase and Quadrature (IQ) data [lehtomaki2005, Juliano2020, Mello2018]. It is actually more common for modern solid state radars to transmit low probability of intercept (LPI) waveforms that utilize LFM pulses at low and negative SNR [Xie2025, pace2009detecting]; in this case, the envelope detector will no longer be suitable. In the low power setting, there are two broad classes of methods, the first operates on the IQ data directly, while the second utilizes time-frequency transforms approaching detection and separation as a 2D signal processing problem. In the time domain, there are analytic methods that operate on the IQ data directly, such as variations of independent components algorithm (ICA) [Joshil2020, Lin2020, Jiang2015, Shao2020], with extensions to nonlinear chirps [PENG2024, zhou2018], and incorporation of simple neural networks [pang2022]. A more common approach is to utilize time-frequency transforms and process the result as an image instead of operating on the IQ data directly [Jin2022, Swiercz2020, Hou2020, Gongming2019, Liu2015, Su2022].
@@
The quantity $\phi_{j}^{\prime}(t)$ is known as the instantaneous frequency (IF) of the component $f_{j}$. The problem is to estimate the IFs and the IAs. This is also a very important and old problem, starting with a paper by a Nobel Laureate, Gabor [gabor1946]. One of the first approaches to solve the problem is the Hilbert-transform based empirical mode decomposition (EMD) by Huang [huang1998empirical]. However, there is no mathematical theory behind this approach. A mathematically rigorous approach was given by Daubeschies and her collaborators [daubechies1996nonlinear, ingrid2011]. These seminal papers have given rise to a lot of variations and improvements. A complete survey is both beyond the scope of this paper as well as not relevant here. We refer to [li2022chirplet] for more references on this topic. In our paper, we will use the algorithm given in [thakur2013synchrosqueezing] for comparing our method with SST.
In [bspaper], we have applied the technique based on localized trigonometric kernels to define a signal separation operator (SSO) to solve this problem under conditions weaker than those in [ingrid2011]. Our conditions do assume that all the components are present in the interval $[0,T]$; i.e., the number of signals $f_{j}$ is the same throughout the interval, and there is no discontinuity in any of these. Our method finds the IFs by looking at the peaks of a power spectrum rather than an elaborate reference curve mechanism of SST. Likewise, the IAs are estimated by a simple substitution compared to the limiting process required in SST. Finally, our method can be implemented very fast using FFT.
The paper [li2022chirplet] deals specifically with the separation of chirplets, based on a modification of a filter described in [bspaper]. It is not clear when the various conditions in this paper are satisfied. Also, the method requires a three dimensional search. In this paper, we will demonstrate how the methods developed in [bspaper] can be adapted for the chirp separation problem, including the case of crossover frequencies.
@@ 3 Mathematical foundations
In this section, we start in Section 3.1 by explaining the basic idea and theorem behind our methods in the case of signals with constant amplitudes and frequencies. In Section 3.2, we explain how the general problem of blind source signal separation can be reduced to this case at each point under certain assumptions as in [ingrid2011, bspaper]. A theoretical analysis of the performance of SSO in the presence of noise is given in Section 3.3.
@@ 3.1 Constant amplitudes and frequencies
where $\epsilon_{\ell}$’s are independent realizations of a sub-Gaussian random variable. This discussion summarizes the corresponding results in [mhaskar2024robust]. Following [boucheron2013concentration, Section 2.3], we recall that a real-valued random variable $X$ is called sub-Gaussian with parameter $V$ ($X\in\mathcal{G}(V)$) if $\mathbb{E}(X)=0$ and $\log\mathbb{E}(\exp(tX))\leq tV^{2}/2$ for all $t>0$; e.g., if $X$ is a mean zero normal random variable with variance $V$, then $X\in\mathcal{G}(V)$. Likewise, any bounded random variable is sub-Gaussian. We define a complex valued random variable $X$ to be in $\mathcal{G}(V)$ if its real and imaginary parts are in $\mathcal{G}(V)$. It is not difficult to see that if $X_{1},\cdots,X_{n}$ are i.i.d., complex valued variables all in $\mathcal{G}(V)$, $\mathbf{a}=(a_{1},\cdots,a_{n})\in{\mathbb{C}}^{n}$, $|\mathbf{a}|_{n}^{2}=\sum_{\ell=1}^{n}|a_{\ell}|^{2}$, then $\sum_{\ell=1}^{n}a_{\ell}X_{\ell}\in\mathcal{G}(|\mathbf{a}|_{n}V)$, and hence, that
@@
Explicit expressions for $L$ in terms of $H$ and $S$ are given in [singdet]. The estimate (3.7) implies that $\hbar_{n}^{-1}\Phi_{n}(x)$ is an approximation to the Dirac delta $\delta_{0}$.
@@
Because of the localization of $\Phi_{n}$, one can deduce that $\sigma_{n}(\mu)(x)$ is small when $x$ is away from each of the $\lambda_{j}$’s, and when it is large then $x$ must be close to exactly one of the $\lambda_{j}$’s. Moreover, $E_{n}(x)$ being a weighted average of the noise realizations is small for large enough $n$. These observations are made precise in the following Theorem 3.1, proved in [mhaskar2024robust].
@@
Before stating the main theorem, we estimate the noise term $E_{n}(x)$ in the following lemma [mhaskar2024robust, Lemma 4.1].
@@ 3.2 Reduction to the constant parameter case
The following theorem is a reformulation of [bspaper, Theorem 4.2]. We will reproduce the proof for the sake of completeness.
@@ 3.3 Signal separation operator
This leads to the following definition (cf. [bspaper, Definition 2.3])
@@
where for each $t$, $\epsilon(t)$ is a realization of a complex sub-Gaussian random variable in $\mathcal{G}(V)$. Theorem 3.1 can be translated into Theorem 3.3 below, where we abuse the notation to make the comparison easier. We note that Theorem 3.3 itself is a restatement of [bspaper, Theorems 2.4,4.1] in a somewhat modified form taking the noise into account. In order to state this theorem, we abuse the notation in Section 3.1 again, and write (in addition to (3.19))
@@ 4 Algorithms
The clusters defined in Theorem 3.3 are found, using a clustering algorithm. For example, in this paper we used DBSCAN [ester1996density]. The notation we use in our algorithm is
@@
Algorithm 1 is designed to find the IFs. A challenge here is to determine the correct threshold related to the unknown minimal amplitude $\mathfrak{m}$ in the presence of noise. A value which is too small would indicate superfluous frequencies, a value which is too large would miss some frequencies. This effect is discussed in our paper [mhaskar2024robust]. There are other details such as detecting the clusters and find the right peaks which are described below. Part 2 will estimate the parameters of the instantaneous frequencies obtained by part 1 using least squares. The challenge here is to figure out where the different signal components begin and end, and which part of the signal to use in order to set up the least square problem in order to obtain reliable solutions. SSO does not work directly when there are crossover frequencies. In fact, the set up itself does not make sense in general in such cases; e.g., whether the crossover of the form $X$ represents two linear frequencies crossing or two $V$-shaped frequencies touching each other. Part 3 is the refinement process where we use the facts that SSO is inherently unable to resolve crossovers and that we are dealing with linear chirps to detect the presence of such frequencies. If sufficiently long linear chirps are present on both sides of the crossover, then we can detect those using the previous two algorithms with a smaller value of $\Delta$ and interpolate to determine where the crossover happens and which chirps cross each other.
@@ 7.3 Comparison with SST
We recall that the most often used method for finding IF’s is EMD, which is purely heuristic. The method known as the synchrosqueezing transform method (SST) is based on a solid mathematical foundation and works better than EMD. In this section, we compare our results with those obtained by using SST. As mentioned earlier, there are many versions of SST. We use the implementation of the SST algorithm as described in Section III of [thakur2013synchrosqueezing].
In the experiment, we use the code from GitHub repository (https://github.com/ebrevdo/synchrosqueezing) from [thakur2013synchrosqueezing]. We use the same parameters used by the authors in their examples, provided in the code, which use the shifted Gaussian (or Morlet) wavelet with zero mean. We refer the reader to [thakur2013synchrosqueezing] for a more detailed discussion of the approach and parameters.
@@ 8 Conclusions
Separation of a superposition of blind source signals is an important problem in audio and radar signal processing, with many applications, including electronic warfare and electronic intelligence. In this paper, we have modified a method proposed in [bspaper] for solving this problem in the case when the instantaneous frequencies of the components are linear. The method is based on a filtered FFT and hence is very fast. While the paper [bspaper] assumes the frequencies to be slowly and continuously varying, our modification includes the case when the signals are discontinuous and when there are crossover frequencies present. We analyze theoretically the behavior of SSO in the presence of noise, giving insight into the relationship between the sampling frequency, desired accuracy, and noise. We have illustrated the effectiveness of our method with a few well chosen examples and tested it on a small database of 7 signals, demonstrating consistent behavior. Our method outperforms the SST algorithm given in [thakur2013synchrosqueezing].